# Copying style, Extracting value: Illustrators’ Perception of AI Style Transfer and its Impact on Creative Labor

### v1

@@ Abstract.
$c$s: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Yet, despite the central role of style in their work, computer scientists have not fully defined the concept or its relation to specific art history approaches or industry realities, highlighting the need for critical examination (we note the seminal exception of (Elgammal et al., 2018)). This lack of specificity is characteristic of the “view from nowhere” approach to technology criticized by feminist scholars (Haraway, 1988), and is potentially harmful to artists. Equating what is generated by style transfer to the rich and complex reality of artists’ style has contributed to consequential narratives of artists’ obsolescence and furthered their alienation from artistic production. Instead, style is best understood specifically, located in the political economy of creative labor. Foregrounding illustrators’ perspectives, this paper asks two questions:
@@
We invited four illustrators to generate images with a diffusion model fine-tuned to their own work to investigate these questions (see Figures 1 and 5). We then conducted semi-structured interviews about their evaluation of the quality, similarity, and usefulness of the output compared to their practice. They were also asked to evaluate the success of the other participants’ models. These results were contextualized in a larger-scale ethnographic fieldwork conducted by one of the authors with the same illustrators.
@@
We contextualize these findings within the political economy of creative labor to show how the design and limitations of style transfer create boundary objects that hinder collaboration with artists but facilitate the extraction of value from their work by their clients. By reproducing a decontextualized, shallow, and predictable copy of style, we argue that rather than a creativity support tool (CST), style transfer is best understood as a supply-chain optimization tool.
@@ 3. Related work
3.1. Theoretical framework
Style sits at the intersection of many stakeholders (researchers, artists, their clients, etc.) who desire to do different things with images. How can we then understand the role of style (and style transfer) in the interactions between these various actors and generative AI? To answer this, we propose conceptualizing style as a “boundary object.”
Understanding style as a boundary object enables us to address the plurality of style perceptions to question the validity of universalizing metrics to evaluate stylistic similarity. This means considering the designers and users of style transfer as partially interacting with style from different perspectives (Velt et al., 2020; John et al., 2004; Jennings, 2005). Engaging with the critical literature that has questioned the universality of datasets (Crawford and Paglen, 2021; Scheuerman et al., 2021; Denton et al., 2021), output (Wevers, 2018), or the positionality of data workers (Miceli and Posada, 2022; Sap et al., 2022), we use boundary objects to show that what tasks like style transfer generate also embed a standpoint (Bardzell, 2010) [^1] While our aim here is to tend to the social plurality of style, we acknowledge there is an intuitive common perception of it grounded in cognition(Augustin et al., 2008; Cupchik et al., 2009).
Boundary objects also raise our awareness of power dynamics in the generation, perception, and control of styles. By facilitating or hindering collaboration between different communities, they can contribute to the power imbalance between them (Chung and Adar, 2023). In the case of the style generated by style transfer, we will show that it hinders collaboration with artists but facilitates the capitalist extraction of value from their work by their clients. In doing so, we link style transfer to the political economy of creative labor, responding to feminist and STS scholars’ call to connect technological solutions to their political context and impact (Winner, 2017; Suchman, 2002; Rogaway, 2015).
@@ 3.2.1. Style as a computer vision challenge
Long before deep neural networks, Aarong Hertzman first proposed “image analogies” (Hertzmann et al., 2001), the first technique for learning generalized style transfer from a pair of transformed images and reproducing an analogous process on a new image. With the advances in deep learning, style transfer dramatically improved, notably with the seminal work of Gatys introduced Neural Style Transfer (Gatys et al., 2016). Using Convolutional Neural Networks, Neural Style Transfer proved to go beyond low-level features of the image to focus on a high-level content-style disentanglement to explicitly extract style from an image and apply it to another (Singh et al., 2021).
@@ 3.2.2. HCI research on style
To our knowledge, no HCI study paper has explicitly focused on artists’ perceptions of style transfer results. We review work on system design, both for and against style mimicry, that has included surveys on artists’ perceptions as well as studies of user interactions with generative AI models in general.
@@
While style transfer is here to stay and has proven useful to some users, it also had a negative impact on artists. A survey of the text-to-image community reveals that (Sanchez, 2023) artist names are the main specifier used alongside the main subject. This has caused much tension between artists and AI companies (Heikkilä, 2022), who in some cases banned the use of living artists’ names as prompts (Vincent, 2022).
Regarding style transfer, the infamous case of a Reddit user who fine-tuned Stable Diffusion to copy illustrator Holly Mengert’s work (Baio, 2022) became a clear example of how illustrators’ styles might be at risk of being simply “grabbed” online to be reproduced by anyone. In response to such threats, Shan et al. (Shan et al., 2023) have developed Glaze, a technique to “cloak” artists’ styles that while imperceptible by the human eye, misleads AI system trying to copy them. Similarly, Liang et al. (Liang et al., 2023) introduced the AdvDM algorithm, which crafts adversarial examples that disrupt the ability of diffusion models to capture and imitate specific artistic features. Meanwhile, Nightshade (Shan et al., 2024) proposes a prompt-specific poisoning strategy named Hemlock, designed to safeguard intellectual properties by subtly destabilizing text-to-image generative models, thus limiting their ability to accurately respond to targeted prompts or rendering them ineffective.
@@ 3.2.3. Style as an underdefined multifaceted object
To understand this semantic diversity of style in the literature and its pragmatic diversity in the creative industry, we introduced the concept of “boundary objects.” This will enable us to consider the interactions between researchers’, illustrators’, and clients’ conceptions of style and their unfolding in the political economy of creative labor.
@@
4. Methodology
4.1. Finetuning text-to-image models
Our study employed Stable Diffusion v1.5 as our base model to personalize text-to-image models. To accomplish this, we curated a dataset consisting of 30 images from each participating illustrator’s online portfolio, which were obtained with their consent. Each image was automatically captioned using BLIP (Li et al., 2022). For the finetuning process, we utilized LoRA (Hu et al., 2021), which is a popular technique known for its ability to adapt models efficiently with minimal additional parameters. The finetuning process executes a total of 3000 training steps with a learning rate of 5e-4 for each participant’s dataset. The training was conducted using Colab Pro GPU, and models were securely stored locally. Initial validation of the training workflow was performed using data from the first author, who is also a proficient illustrator. Each participant’s model was then made accessible for experimentation via the stable diffusion web interface [^2] https://github.com/AUTOMATIC1111/stable-diffusion-webui(Figure 2).
@@ 4.2.1. Participants
We recruited four artists (2 females and 2 males), all professional illustrators at the moment of the study. Participants’ ages ranged from 29 to 33 years old, and they all had been working as illustrators between 8 and 11 years. Each illustrator has worked with editorial, advertising, and/or branding clients in the US and internationally. The recruitment was conducted through contacts established by one of the authors during his long-term fieldwork in the illustration industry. Our main eligibility criteria was working professionally as an illustrator, part-time or full-time. This was an essential choice as we understand style holistically, contributing to artistic and economic value. Therefore, the insights of these professional illustrators on the value of their style in both contexts were crucial. This choice also contextualizes artistic practice (and style) in broader social systems instead of upholding the cultural narrative of art’s exceptionalism (Becker, 2023).
Participants were also selected based on the diversity of their styles and mediums to enrich the diversity of perception we are looking to make salient. None of the participants had extensive experience with generative AI and were generally curious about the perspective of having a model trained on their work. It is important to note that this was not the case for all the illustrators approached for this study. Some illustrators categorically refused to participate, stating, for example, that the perspective of seeing their lifework reproduced by a machine would be too challenging. Participants’ explicit consent was obtained after a thorough explanation of the study procedure as well as the ethical measures taken to protect them.
@@ 4.2.2. Study Procedure
Participants were given access to their fine-tuned models via a stable-diffusion app to generate images during sessions that ranged from two to three hours. Participants were free to prompt the model however they wished, with the exception of a ”control prompt” common to all: ”the last supper” (Figure 6). Semi-structured interviews took place during and after the study, adapting to the participants’ reactions to the generated output in the moment and exploring their overall perception at the end. Interviews focused on their perception of the quality of the output in general and as a copy of their own style. They were also asked to reflect on usability and past professional experiences related to style transfer. Finally, each participant was shown other participants’ model’s output and was asked to evaluate its success in copying their style.
@@ 5. Findings
Our findings show that evaluating style transfer’s success, far from being a neutral assessment of a model’s performance, is a socially situated perceptual task, i.e., not everyone sees the output of style transfer the same way. Our questions focused on artists’ perception of (1) what was successfully reproduced of their style, (2) what made the model unable to reproduce their style, (3) the usefulness of the model in their work, and (4) their perception of other participants’ output. These themes will outline significant differences in the ways AI models and artists perform style.
@@ 5.1. Texture, colors, and other successful fragments
When the participants were asked to evaluate the model’s success in copying their style, they identified isolated elements of the images, referencing low-level aspects of the images such as colors, textures, shading, and so on. At first, struggling to find anything successful in this copy, closer inspection would lead them to identify small pieces of the images they saw as accurate depictions of their style, stating that “[…] some of these underarm shadows are pretty good” [P3] or “the lines on the shirt, I think it picked up on that. When I draw shirts, I would do folds and stuff [P4].” This granular evaluation of stylistic output contrasted usual evaluations of style transfer which tends to deal with “style” broadly defined, if defined at all. We now turn to texture, color, and shading in particular.
@@
5.1.1. Texture
Texture was generally successfully transferred from artists’ work to the output, for example, evaluating another artist’s [P2] output, this illustrator was impressed by its quality:
@@ 5.1.1. Texture
Another commented on a specific texture coming from a particular image from the dataset:
[P2]: “Well, better than anyone I just see specific little things that I feel like that might have been just taken from one of the images, and that’s like where the texture comes from”
@@ 5.1.2. Color
Another aspect of the output that artists positively evaluated was color. Most of them were surprised and pleased by the accuracy of the color palette generated in the output, stating “Then you see the composition and you see the colors and I think that’s what it does best” [P2] and “The colours are quite successful. That’s an interesting thing” [P1]. Even in assessing other participants, color was a key successful element - [P1]: “I think color wise they nailed her [P2]”. Recognizing someone’s work by their color palette points to how important color is to style as an index of their identity and uniqueness.
@@ 5.1.3. Shading
Shading is obviously limited to particular elements of the images and was not always present in the output (Figure 4). Yet, even when the overall results did not convince them, participants often found the shading to be successful.
Figure 4. Shading evaluated as successfully copied
[…]
@@ 5.1.4. Textures, colors, and shading did not equate style
Textures, colors, and shading are all discreet elements whose results never exceeded the sum of their parts for the artists involved. The model successfully reproduced such isolated, low-level elements of each dataset but only in localized places and not equally throughout every output. Most importantly, artists’ evaluations show that however successful the model was at reproducing some parts of their style, it was only doing so in a way that never exceeded the sum of its parts:
@@
[P2]: “I mean, the same way I can see me in my images, like they are compositional things, like the kind of relation of how big is a head compared to a body, the color choices and some little artifacts like mark-making, it gets right, but then it just chops them up so weirdly”
This “chopped-up” version of style mimicry points to a crucial difference in how artists and models perform style. Interestingly, our findings closely match the features Garces et al. (Garces et al., 2014) have proposed as metrics for stylistic similarity: texture, shading, colors, and stroke. Yet, as we’ve shown, while the artists deemed these elements of style transfer successful, it was always in a narrow and limited sense. In other words, while the model successfully transferred some qualities of the dataset, these could hardly be qualified to amount to the artists’ style.
@@ 5.2. Challenging the aesthetic-content disentanglement task in style transfer
When asked about the model’s limitations in generating a successful copy of their style, participants often pointed to the difficulty of working with style-content disentanglement.
@@ 5.2.2. Semantic objects are stylistically important
Each artist has semantic elements they always draw, or conversely never do, the presence and absence of which become an essential part of their style over time. Many participants prompted the models to generate subjects they generally never draw, and therefore absent from the dataset. In such instances, the model struggled to generate output that was interesting to them. As this illustrator explains:
@@
The result of the prompt “a zoo and a circus” generated an output considered an outlier compared to the more successful ones that participants generated. She considered the result very unsuccessful and joked that she too would actually be unsuccessful at drawing animals:
@@
Semantic elements therefore contribute to style as a whole, rather than be separate from it.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713854isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
[…]
4 pairs of 8 illustrations are presented on a grid. Each pair comprises four illustrations created by one of the participants on the left-hand side and four illustrations generated by our model on the right-hand side. Above the illustrator’s four images is their participant code (Pn), and above the generated images is Pn’. Each pair gives an overview of general stylistic similarity between the illustrators’ work and their models.
@@
Yet, despite the central role of style in their work, computer scientists have not fully defined the concept or its relation to specific art history approaches or industry realities, highlighting the need for critical examination (we note the seminal exception of (Elgammal et al., 2018)). This lack of specificity is characteristic of the “view from nowhere” approach to technology criticized by feminist scholars (Haraway, 1988), and is potentially harmful to artists. Equating what is generated by style transfer to the rich and complex reality of artists’ style has contributed to consequential narratives of artists’ obsolescence and furthered their alienation from artistic production. Instead, style is best understood specifically, located in the political economy of illustration as a profession. Foregrounding illustrators’ perspectives, this paper asks two questions:
@@
We invited four illustrators to generate images with a diffusion model fine-tuned to their own work to investigate these questions (see Figures 1 and 7). We then conducted semi-structured interviews about their evaluation of the quality, similarity, and usefulness of the output compared to their practice. They were also asked to evaluate the success of the other participants’ models. These results were contextualized in a larger-scale ethnographic fieldwork conducted by one of the authors with the same illustrators.
@@
We contextualize these findings within the political economy of illustration work to show how the design and limitations of style transfer create boundary objects that hinder collaboration with artists but facilitate the extraction of value from their work by their clients. By reproducing a decontextualized, shallow, and predictable copy of style, we argue that rather than a creativity support tool (CST), style transfer is best understood as a supply-chain optimization tool.
@@ 3.1.1. Style as a computer vision challenge
Long before deep neural networks, Aaron Hertzman first proposed “image analogies” (Hertzmann et al., 2001), the first technique for learning generalized style transfer from a pair of transformed images and reproducing an analogous process on a new image. With the advances in deep learning, style transfer dramatically improved, notably with the seminal work of Gatys introduced Neural Style Transfer (Gatys et al., 2016). Using Convolutional Neural Networks, Neural Style Transfer proved to go beyond low-level features of the image to focus on a high-level content-style disentanglement to explicitly extract style from an image and apply it to another (Singh et al., 2021).
@@ 3.1.2. HCI research on style
To our knowledge, no HCI study paper has explicitly focused on artists’ perceptions of style transfer results. We review work on system design, both for and against style mimicry, that has included surveys on artists’ perceptions and studies of user interactions with generative AI models in general.
@@
While style transfer is here to stay and has proven useful to some users, it also had a negative impact on artists. A survey of the text-to-image community reveals that (Sanchez, 2023) artist names are the main specifier used alongside the main subject. This has caused much tension between artists and AI companies (Heikkilä, 2022), who in some cases banned the use of living artists’ names as prompts (Vincent, 2022). Furthermore, in (Srinivasan and Uchino, 2021), Srinivasan et al. show that modeling artists’ styles using correlation statistics often leads to stereotyping the nature of their process and the quality of their work, leading to a biased representation of what style is.
Regarding style transfer, the infamous case of a Reddit user who fine-tuned Stable Diffusion to copy illustrator Holly Mengert’s work (Baio, 2022) became a clear example of how illustrators’ styles might be at risk of being simply “grabbed” online to be reproduced by anyone. In response to such threats, Shan et al. (Shan et al., 2023) have developed Glaze, a technique to “cloak” artists’ styles that while imperceptible by the human eye, misleads AI system trying to copy them. Similarly, Liang et al. (Liang et al., 2023) introduced the AdvDM algorithm, which crafts adversarial examples that disrupt the ability of diffusion models to capture and imitate specific artistic features. Meanwhile, Nightshade (Shan et al., 2024) proposes a prompt-specific poisoning strategy, designed to safeguard intellectual properties by subtly destabilizing text-to-image generative models, thus limiting their ability to accurately respond to targeted prompts or rendering them ineffective.
@@ 3.1.3. Style as an underdefined multifaceted object
4. Methodology
4.1. Finetuning text-to-image models
@@
Screenshot of the Stable Diffusion Web User Interface the participants used for the study. The page consists of a prompt bar in which is written the prompt “a still life of a vase, a fruit and a cup on a table”; below it a negative prompt bar, and to their right, a big orange “Generate” button. Below the prompt text bars are a series of sliders for image width, height, batch count, batch size, and sampling steps. To the right of these sliders, under the orange button, is the generated image of a flower vase on a table with some fruits.
Our study employed Stable Diffusion v1.5 as our base model to personalize text-to-image models. To accomplish this, we curated a dataset consisting of 30 images from each participating illustrator’s online portfolio, which were obtained with their consent. Each image was automatically captioned using BLIP (Li et al., 2022). For the finetuning process, we utilized LoRA (Hu et al., 2021), which is a popular technique known for its ability to adapt models efficiently with minimal additional parameters. The finetuning process executes a total of 3000 training steps with a learning rate of 5e-4 for each participant’s dataset. The training was conducted using Colab Pro GPU, and models were stored locally in a secure and password-encrypted device. Initial validation of the training workflow was performed using data from the first author, who is also a proficient illustrator. Each participant’s model was then made accessible for experimentation via the stable diffusion web interface [^1] https://github.com/AUTOMATIC1111/stable-diffusion-webui(Figure 2).
@@ 4.2.1. Participants
We recruited four artists (2 females and 2 males, background see Table 1), all professional illustrators working in New York City at the moment of the study. New York City is a nodal point for the illustration industry, with many publications and agencies located there. The city, with its high cost of living, is also a place where illustrators are particularly aware of the political economy of their craft, which made their insights on the matter all the more vivid. Participants ranged from 29 to 33 years old, and they all had been working as illustrators between 8 and 11 years. Each illustrator has worked with editorial, advertising, and/or branding clients in the US and internationally. The recruitment was conducted through contacts established by one of the authors during his long-term fieldwork in the illustration industry. Our main eligibility criteria was working professionally as an illustrator, part-time or full-time. This was an essential choice as we understand style holistically, contributing to artistic and economic value. Therefore, the insights of these professional illustrators on the value of their style in both contexts were crucial. This choice also contextualizes artistic practice (and style) in broader social systems instead of upholding the cultural narrative of art’s exceptionalism (Becker, 2023).
Participants were also selected based on the diversity of their styles and mediums to enrich the diversity of perception we are looking to make salient. None of the participants had extensive experience with generative AI and were generally curious about the perspective of having a model trained on their work. It is important to note that this was not the case for all the illustrators approached for this study.
Explicit consent was obtained after conversations with the participants over months prior to the study. Embedded in the social network of these illustrators for a year of ethnographic fieldwork, the first author discussed the idea of this study first informally during conversations about AI and illustration, then more formally with each participant, thoroughly explaining the study procedure as well as the ethical measures taken to protect them. As we have discussed, illustrators’ relationship to AI is one of dispossession, their work being fed to systems without their consent. Seven illustrators were approached, and three refused. One of them refused because they were not interested in the topic, and two because they felt that seeing their lifework reproduced easily by a machine would be too confronting. Refusal to participate in a study is a social position towards technological systems that is rich in insights, articulating, for example, a fear of exploitation (Grigis and De Angeli, 2024). In this context, the choice of having such a small sample has allowed us to develop trusting relationships with each participant. These relationships are not only about obtaining consent from participants, but also creating accountability for researchers. We return to the topic of participant selection in section 7.
@@ 4.2.2. Study procedure
Participants were given access to their fine-tuned models via a stable-diffusion app to generate images during sessions that ranged from two to three hours. Participants were free to prompt the model however they wished, with the exception of a “control prompt” common to all: “the last supper” (Figure 6). Semi-structured interviews took place during and after the study, adapting to the participants’ reactions to the generated output in the moment and exploring their overall perception at the end. Interviews focused on their perception of the quality of the output in general and as a copy of their own style. They were also asked to reflect on usability and past professional experiences related to style transfer. Finally, each participant was shown other participants’ model’s output and was asked to evaluate its ability to copy their style. Using an ”emergent coding” approach, i.e., inductive coding (Lazar et al., 2017), we conducted a thematic analysis (Braun and Clarke, 2006) of the interview transcripts using the software Atlas.ti to note recurring patterns in the participants’ responses. After the first coding pass, the team discussed the various codes to come to an agreement on the most important ones. The high-level categories have been used to structure the Findings section to stay as close as possible to the data.
@@ 5. Findings
Our findings show that evaluating style transfer’s success, far from being a neutral assessment of a model’s performance, is a socially situated perceptual task, i.e., not everyone sees the output of style transfer the same way. Our questions focused on artists’ perception of (1) what was accurately reproduced of their style, (2) what made the model unable to reproduce their style, (3) the usefulness of the model in their work, and (4) their perception of other participants’ output. These themes will outline significant differences in the ways AI models and artists perform style.
@@ 5.1. Texture, colors, and other successful fragments
When the participants were asked to evaluate the model’s success in copying their style, they identified isolated elements of the images, referencing low-level aspects of the images such as colors, textures, shading, and so on. At first, struggling to find anything similar to their work in this copy, closer inspection would lead them to identify small pieces of the images they saw as accurate depictions of their style, stating that “[…] some of these underarm shadows are pretty good” [P3] or “the lines on the shirt, I think it picked up on that. When I draw shirts, I would do folds and stuff [P4].” This granular evaluation of stylistic output contrasted usual evaluations of style transfer which tends to deal with “style” broadly defined, if defined at all. We now turn to texture, color, and shading in particular.
5.1.1. Texture
@@
2 pairs of 2 images are stacked on each other; the top pair is labeled P4’, and the bottom is labeled P2’. In each pair, the first image is an AI generated image with a white, 80% opacity coverage, leaving the image barely visible except for one small square left clear. The second image of each pair is that small clear square of the first image blown up, showing the texture of a small patch of an image.
Texture was generally accurately transferred from artists’ work to the output, for example, evaluating another artist’s [P2] output, this illustrator was impressed by its quality:
@@
Here, the texture P4 references is not a quality of the entire image but a localized element in the generation of clouds only. Another commented on a specific texture coming from a particular image from the dataset:
[P2]: “Well, better than anyone I just see specific little things that I feel like that might have been just taken from one of the images, and that’s like where the texture comes from.”
@@ 5.1.2. Color
Another aspect of the output that artists positively evaluated was color. Most of them were surprised and pleased by the accuracy of the color palette generated in the output, stating, “Then you see the composition and you see the colors and I think that’s what it does best” [P2] and “The colors are quite successful. That’s an interesting thing” [P1]. Even in assessing other participants, color was an accurately copied element: “I think color wise they nailed her” says [P1], talking about what the model succeeded in reproducing about [P2]’s work. Recognizing someone’s work by their color palette points to how important color is to style as an index of their identity and uniqueness.
@@ 5.1.3. Shading
Figure 4. Shading evaluated as successfully copied
[…]
A pair of images labeled P2’. The first image is an AI generated image with a white, 80% opacity coverage, leaving the image barely visible except for one small square left clear. The second image is that small clear square of the first image blown up, showing the shading under a character’s dress.
@@
Shading is obviously limited to particular elements of the images and was not always present in the output (Figure 4). Yet, even when the overall results did not convince them, participants often found the shading to be accurately copied.
@@ 5.1.4. Textures, colors, and shading did not equate style
Textures, colors, and shading are all discreet elements whose results never exceeded the sum of their parts for the artists involved. The model was good at reproducing such isolated, low-level elements of each dataset but only in localized places and not equally throughout every output. Most importantly, artists’ evaluations show that however successful the model was at reproducing some parts of their style, it was only doing so in a way that never exceeded the sum of its parts:
@@
[P2]: “I mean, the same way I can see me in my images, like they are compositional things, like the kind of relation of how big is a head compared to a body, the color choices and some little artifacts like mark-making, it gets right, but then it just chops them up so weirdly.”
This “chopped-up” version of style mimicry points to a crucial difference in how artists and models perform style. Interestingly, our findings closely match the features Garces et al. (Garces et al., 2014) have proposed as metrics for stylistic similarity: texture, shading, colors, and stroke. Yet, as we’ve shown, while the artists deemed these elements of style transfer indistinguishable from their work, it was always in a narrow and limited sense. In other words, while the model successfully transferred some qualities of the dataset, these could hardly be qualified to amount to the artists’ style.
@@ 5.2. Challenging the aesthetic-content disentanglement task in style transfer
When asked about the model’s limitations in generating an accurate copy of their style, participants often pointed to the difficulty of working with style-content disentanglement.
@@ 5.2.2. Semantic objects are stylistically important
Figure 5. Two style transfer outputs from the prompt ”Zoo animals and circus” by [P2]
[…]
Two illustrations side by side. The illustrations were generated by P2 using our style transfer model with the prompt ”Zoo animals and circus.”
Each artist has semantic elements they always draw, or conversely never do, the presence and absence of which become an essential part of their style over time. Many participants prompted the models to generate subjects they generally never draw, and therefore absent from the dataset. In such instances, the model struggled to generate output that was interesting to them.
As this illustrator explains:
@@
[P1] expresses here the difficulty of picturing her style, including the semantic elements of animals. Thus the result of the prompt “a zoo and a circus” generated an output considered an outlier compared to the more accurate copies that participants generated (Figure 5). She considered the result not to be representative of her style and joked that she too would actually be bad at drawing animals:
@@
As [P2] expresses here, even within such a semantically normative genre as the self-portrait, what made her style hers was the alteration of the ”subject of the content of the illustration” alongside the aesthetic aspect of it.
Figure 6. Images generated with the control prompt ”the last supper”
[…]
A series of 4 illustrations side by side, each labeled from P1’ to P4’. They all depict the Last Supper in what the model has interpreted as the style of each participant.
Prior to the interviews, we generated some images with each participant’s model using generic prompts such as ”two persons talking” or the control prompt ”the last supper” (see Figure 6). Some participants found that the model reproduced their style less accurately with these prompts than their own. For example, some of P2’s prompts were:
”a tree and and over proportionally big flower on an island in the foreground with a person sleeping underneath it. in the background a lake and mountain. wind is blowing through the tree. it is dusk. birds are flying.”
or
”a man is standing in an empty street, reaching up to touch a bird is flying above him.”
Similarly, P3 spent a considerable amount of time prompting an image of a character he enjoys drawing, Humpty Dumpty, iterating numerous times to give the right stylistic cues:
”An egg person is sitting on a brick wall. He is an egg with a face and short arms and legs. His hands are very big. His world is not three dimensional. He has a big blank smile. His eyes are blank. He wears a bowtie. His clothes are simple, they do not have intricate features. There is an apple sitting next to him on the wall.”
These prompts, unlike our control ones, are rich with idiosyncratic semantic elements (P2’s birds, over proportionate flowers, or P3’s Humpty Dumpty with a bowtie) and stylistic directions, which are intrinsic to their evaluation of style transfer as accurate. Commenting on the difference between the output generated by our respective prompts, P2 explains:
[P2]: “It takes me a lot of time and effort to see it [the similarity]. Especially with these images that you generated. The other ones [that she generated during the experiment] that was quicker to see the resemblances because those are also things that are different from my work and different from how I approach drawing.”
Evaluating the ability of the model to reproduce style was highly dependent on the semantic elements chosen to generate output. This points once again to the impossibility to disentangle style from content, or at least its uselessness for artists. Semantic elements therefore contribute to style as a whole, rather than be separate from it.